Tanner Corporation produced 4,270 units, consisting of three separate products, in a joint process for the year. The market for these products was so unstable that it was not practical to estimate the selling price of the products. A cost of $471,000 was incurred in the joint process. Product X's production was 80% of product Y's while product Z's production was 125% of product Y's. What is the amount of the joint cost allocable to product X (rounded to the nearest whole dollar) assuming Tanner uses the physical quantities method of allocation?
Answer #1

Lets product Y production is equal to 100%

Then Product X = 100*80% = 80%

Product Z = 100*125% = 125%

Product Percentage Physical Unit Joint Cost Joint cost
X 80% 4270*80/305 = 1120 471000 471000*1120/4270 = 123541
Y 100% 1400 471000 154426
Z 125% 1750 471000 193033
Total 305% 4270 471000

So answer is $123541